OpenStudy (anonymous):

Series & Sequences: Each year a school manages to use only 90% as much paper as the previous year in the year 2000, they used 700,000 sheets of paper. a) Find how much paper the school used in the years 2001 and 2002. b) How much total paper did the school use in the decade from 2000 to 2009?

OpenStudy (anonymous):

Do you have any idea of a answer?

OpenStudy (anonymous):

for 2001 it would be \[700,000\times .9\]

OpenStudy (anonymous):

for 2002 it would be \[700,000\times .9\times .9=700,000\times (.9)^2\]

OpenStudy (anonymous):

oh you are asked for the total, sorry then you would have to add \[700,000\times .9+700,000\times (.9)^2\]

OpenStudy (agent0smith):

For (b) you can use the geometric sequence sum formula.